OpenStudy (anonymous):

log6 (x+4) - log6 (x-2)= log6 (x)

OpenStudy (anonymous):

\[\log_{6} (x+4) - \log_{6} (x-2= \log_{6} (x+4)/(x-2) = \log_{6} x\] then take exponential (base 6 ofcource) bouth sides and solve equation

OpenStudy (anonymous):

i dont understand how you get x from that

OpenStudy (anonymous):

taking exponential \[6^{\log 6 (x+4)/(x−2)}= 6^{\log 6 x}\] (x+4)/(x-2) =x x+4 = x(x-2) and solve

OpenStudy (anonymous):

you should get x=4 and x=-1

OpenStudy (anonymous):

how does x(x-2) equal out to -1?

OpenStudy (anonymous):

\[x^{2} -3x -4 =0\]

OpenStudy (anonymous):

ohh. alright thanks. i understand now

OpenStudy (anonymous):

this is what u get if u pass all terms to one side x+4 = x(x-2)

OpenStudy (anonymous):

thank you, can you help me with some other problems on my profile?

Directrix (directrix):

log6 (x + 4) - log6(x-2) = log6(x) log6 [ (x + 4) / (x + 2) ] = log6(x) [ (x + 4) / (x + 2) ] = x x^2 - 2x = x + 4 x^2 - 3x - 4 = 0 (x - 4) (x + 1) = 0 x = 4 or x = -1 x = 4 (x = -1 is an extraneous solution)